# Model Already Knows the Best Noise: Bayesian Active Noise Selection via Attention in Video Diffusion Model

### v1

@@ Appendix C Proof of Proposition 1
Appendix D Further Details on Evaluation Metrics and Implementation
@@
Total $\text{Total Score}=\frac{w_{1}}{w_{1}+w_{2}}\cdot\text{Quality Score}+\frac{w_{2}}{w_{1}+w_{2}}\cdot\text{Semantic Score}$
@@
As discussed in Section B, we compute our BANSA score using the BALD-style formulation, which yields non-negative values. For clearer visualization, we normalize the BANSA scores from their original range (minimum: 0.45, maximum: 0.60) to the [0, 1] interval. This normalization is used solely for visual clarity in figures and plots, and does not affect the noise selection process, which operates on the raw BANSA scores.
Appendix E Further Detail of BANSA Layer-wise Correlation Analysis
@@
For each prompt, we generated 10 videos using different random noise seeds and computed BANSA scores at each attention layer. This yielded one full-layer BANSA score and a set of layer-wise scores per seed.
To obtain stable estimates, we averaged the per-layer and full-layer BANSA scores across the 10 seeds, reducing noise-specific variance and capturing consistent uncertainty patterns.
We then computed Pearson correlations between the cumulative BANSA scores (summed from layer 1 to $d$) and the official quality scores. The optimal depth $d^{*}$ was defined as the smallest $d$ at which the correlation exceeded 0.7. This procedure is visualized in Figure 5 of the main paper, and $d^{*}$ was applied consistently throughout all experiments. This setup ensures that our correlation analysis reflects generalizable, noise-agnostic trends in attention-based uncertainty.
@@
CogvideoX-5B Full-layer 0.9660 0.9630 0.9863 0.9708 0.6168 0.6290 0.6979 82.71 1530.1
Truncated 0.9658 0̱.9639 0.9861 0.9711 0.6179 0.6290 0.6918 82.70 1392.3
Appendix F Effectiveness of Truncated BANSA Score
@@
Appendix G Effect of Temporal Scope in BANSA Score
@@
Appendix H Additional Qualitative Comparison
@@
Figures 8 and 9 present additional examples generated using our noise selection framework. Across diverse prompts, the selected seeds yield improved spatial detail, aesthetic quality, and semantic alignment, further validating the robustness of our approach. These examples complement our quantitative findings by illustrating the visual impact of BANSA-based noise selection.
@@
Figure 8: Effect of ANSE on semantic fidelity and motion stability in CogVideoX outputs. Each block compares baseline generations with those using ANSE-selected noise. Across both CogVideoX-2B and 5B, ANSE improves semantic alignment to the prompt and reduces artifacts such as temporal flickering and object distortion.
Figure 9: Additional qualitative comparison of CogVideoX variants with and without ANSE. Results from CogVideoX-2B are shown in the first two rows; the rest show CogVideoX-5B. With ANSE, videos exhibit improved visual quality, better text alignment, and smoother motion transitions compared to the baseline.

### v2

@@ Appendix B Proof of Proposition 1
Appendix C Further Details on Evaluation Metrics and Implementation
@@
$\text{Total Score}=\frac{w_{1}}{w_{1}+w_{2}}\cdot\text{Quality Score}+\frac{w_{2}}{w_{1}+w_{2}}\cdot\text{Semantic Score}$
@@
We compute our BANSA score using the BALD-style formulation, which yields non-negative values. For clearer visualization, we normalize the BANSA scores from their original range (minimum: 0.45, maximum: 0.60) to the [0, 1] interval. This normalization is used solely for visual clarity in figures and plots, and does not affect the noise selection process, which operates on the raw BANSA scores. All models are evaluated using their default inference configurations: AnimateDiff (50 steps), CogVideo-X (50 steps), Wan2.1 (50 steps), and HunyuanVideo (50 steps).
Appendix D Further Detail of BANSA Layer-wise Correlation Analysis
@@
For each prompt, we generated 10 videos using different random noise seeds and computed BANSA scores at each attention layer. This yielded one full-layer BANSA score and a set of layer-wise scores per seed. To obtain stable estimates, we averaged the per-layer and full-layer BANSA scores across the 10 seeds, reducing noise-specific variance and capturing consistent uncertainty patterns.
We then computed Pearson correlations between the cumulative BANSA scores (summed from layer 1 to $d$) and the official quality scores. The optimal depth $d^{*}$ was defined as the smallest $d$ at which the correlation exceeded 0.7, a widely used threshold indicating a strong positive relationship. As shown in Figure 7, this point was reached at $d^{*}=14$ for CogVideoX-2B and $d^{*}=19$ for CogVideoX-5B. For AnimateDiff, the correlation crossed the threshold earlier at $d^{*}=10$, while for HunyuanVideo and Wanx 2.1 it consistently appeared around $d^{*}=20$. These model-specific depths were applied throughout all experiments. Such consistency across architectures empirically supports our choice of the 0.7 threshold and ensures that the correlation analysis reflects generalizable, noise-agnostic patterns of attention-based uncertainty.
@@
CogvideoX-5B Full-layer 0.9660 0.9630 0.9863 0.9708 0.6168 0.6290 0.6979 82.71 810.3
Truncated 0.9658 0̱.9639 0.9861 0.9711 0.6179 0.6290 0.6918 82.70 754.3
Appendix E Effectiveness of Truncated BANSA Score
@@
Appendix F Effect of Temporal Scope in BANSA Score
@@
Table 14: Quantitative results on VBench using AnimateDiff, CogVideoX-2B and -5B.
Backbone Model Method Quality Score Semantic Score Total Score
AnimateDiff (Guo et al., 2024b) Vanilla 80.22 69.03 77.98
+ Ours (Self-Attention) 81.68 70.72 79.19
+ Ours (Cross-Attention) 81.66 71.09 79.33
Appendix G Effect of Attention Space in BANSA Score
To further examine whether BANSA scores capture quality- or semantic-related uncertainty, we conducted additional experiments with the AnimateDiff backbone, which separates self- and cross-attention layers. Unlike CogVideo and Wanx, which adopt an MMDiT architecture where multimodal attention entangles both components, AnimateDiff provides a clearer lens for disentangling semantic effects. As shown in Table 14, applying BANSA to self-attention primarily improved perceptual quality scores, whereas applying it to cross-attention enhanced semantic alignment. This indicates that uncertainty over noise manifests differently depending on the attention type, supporting the interpretation that BANSA reflects both quality- and semantics-related variations. In contrast, MMDiT-based models such as CogVideo and Wanx integrate these dimensions within their unified attention structure, making them more suitable for overall noise selection in the main paper’s experiments.
Table 15: Ablation study on the interaction between higher CFG scale and ours on Wan2.1(1.3B).
Backbone Model Method CFG Scale Subject Consistency $\uparrow$ Motion Smoothness $\uparrow$ Aesthetic Quality $\uparrow$ Imaging Quality $\uparrow$
Wanx 2.1 1.3B Vanilla 5.0 0.9285 0.9748 0.6465 0.6519
Ours 5.0 0.9310 0.9786 0.6559 0.6516
Vanilla 7.5 0.9266 0.9740 0.6447 0.6357
Ours 7.5 0.9276 0.9755 0.6519 0.6409
Appendix H Effect of CFG Scale on Our Method
Classifier-free guidance (CFG) modifies the conditional–unconditional interpolation only at the final stage of inference, whereas our method selects noise seeds at initialization based on attention-level uncertainty. Since they operate at different points in the pipeline, their effects are largely independent: CFG adjusts the extrapolation strength, while our method chooses seeds that lead to more stable early-stage behavior.
To assess potential interactions, we increased the CFG scale from the default value of 5.0 to 7.5 on the Wanx 2.1 (1.3B) model, as shown in Table 15. Higher CFG values generally degrade performance by pushing sampling beyond the model’s calibrated range. Nevertheless, our method consistently outperforms the vanilla baseline under both settings, indicating that its benefits remain stable across different CFG scales.
Table 16: Ablation study under an equivalent computation budget. Increasing the baseline sampling steps does not consistently improve performance, whereas our method achieves better results with fewer effective steps.
Backbone Model Method Steps Subject Consistency $\uparrow$ Motion Smoothness $\uparrow$ Aesthetic Quality $\uparrow$ Imaging Quality $\uparrow$
Wan2.1 1.3B Vanilla 50 0.9285 0.9748 0.6465 0.6519
Vanilla 60 0.9274 0.9745 0.6476 0.6492
Ours << 60 0.9310 0.9786 0.6559 0.6516
Appendix I Effect of Equivalent Computation Budget
Our method evaluates $M=10$ noise candidates, but the actual computation is lower than $50+10$ steps because many candidates terminate early during partial denoising. To assess whether the baseline could benefit from additional computation, we increased its sampling steps from 50 to 60. As shown in Table 16, this results in only marginal changes and even degrades some metrics, likely because the model is not calibrated for longer sampling schedules.
In contrast, our method consistently improves performance while requiring fewer effective steps than the 60-step baseline, suggesting that the gains stem from selecting more stable noise initializations rather than from additional compute. This highlights the efficiency advantage of our approach under equivalent or lower computation budgets.
[…]
Backbone Model Method Subject Consistency $\uparrow$ (CI) Background Consistency $\uparrow$ (CI) Temporal Flickering $\uparrow$ (CI) Motion Smoothness $\uparrow$ (CI) Aesthetic Quality $\uparrow$ (CI) Imaging Quality $\uparrow$ (CI) Dynamic Degree $\uparrow$ (CI)
CogvideoX-2B Baseline 0.9616 (0.9606 - 0.9626) 0.9788 (0.9779 - 0.9797) 0.9715 (0.9688 - 0.9742) 0.9738 (0.9733 - 0.9753) 0.6195 (0.6116 - 0.6274) 0.6267 (0.6160 - 0.6374) 0.6380 (0.6133 - 0.6627)
+ Ours 0.9639 (0.9629 - 0.9649) 0.9810 (0.9795 - 0.9825) 0.9801 (0.9781 - 0.9821) 0.9743 (0.9737 - 0.9756) 0.6198 (0.6119 - 0.6277) 0.6244 (0.6137 - 0.6351) 0.6500 (0.6255 - 0.6745)
CogvideoX-5B Baseline 0.9616 (0.9608 - 0.9624) 0.9616 (0.9602 - 0.9630) 0.9862 (0.9833 - 0.9891) 0.9712 (0.9703 - 0.9721) 0.6162 (0.6084 - 0.6240) 0.6272 (0.6170 - 0.6374) 0.6895 (0.6655 - 0.7135)
+ Ours 0.9660 (0.9653 - 0.9667) 0.9630 (0.9616 - 0.9644) 0.9863 (0.9835 - 0.9891) 0.9708 (0.9700 - 0.9716) 0.6168 (0.6089 - 0.6247) 0.6290 (0.6189 - 0.6391) 0.6979 (0.6738 - 0.7220)
[…]
CogvideoX-2B Baseline 0.8522 (0.8511 - 0.8533) 0.6391 (0.6376 - 0.6406) 0.8318 (0.8302 - 0.8334) 0.7148 (0.6963 - 0.7333) 0.4815 (0.4801 - 0.4829) 0.2488 (0.2453 - 0.2523) 0.2722 (0.2679 - 0.2765) 0.9860 (0.9743 - 0.9977) 0.2497 (0.2483 - 0.2511)
+ Ours 0.8842 (0.8831 - 0.8853) 0.6596 (0.6581 - 0.6611) 0.8255 (0.8240 - 0.8270) 0.6915 (0.6728 - 0.7102) 0.5286 (0.5272 - 0.5300) 0.2547 (0.2511 - 0.2583) 0.2760 (0.2718 - 0.2802) 0.9870 (0.9757 - 0.9963) 0.2507 (0.2493 - 0.2521)
CogvideoX-5B Baseline 0.8595 (0.8585 - 0.8605) 0.6511 (0.6496 - 0.6526) 0.8218 (0.8202 - 0.8234) 0.6872 (0.6675 - 0.7069) 0.5233 (0.5219 - 0.5247) 0.2552 (0.2513 - 0.2591) 0.2761 (0.2717 - 0.2805) 0.9900 (0.9813 - 0.9987) 0.2483 (0.2354 - 0.2612)
+ Ours 0.8675 (0.8665 - 0.8685) 0.6658 (0.6643 - 0.6673) 0.8201 (0.8186 - 0.8216) 0.6831 (0.6637 - 0.7025) 0.5255 (0.5241 - 0.5269) 0.2585 (0.2546 - 0.2624) 0.2773 (0.2730 - 0.2816) 0.9980 (0.9910 - 1.0050) 0.2524 (0.2511 - 0.2537)
[…]
Backbone Model Method Subject Consistency $\uparrow$ (CI) Background Consistency $\uparrow$ (CI) Motion Smoothness $\uparrow$ (CI) Aesthetic Quality $\uparrow$ (CI) Imaging Quality $\uparrow$ (CI) Temporal Flickering $\uparrow$ (CI)
Hunyuan Video Vanilla 0.9562 (0.9551 – 0.9573) 0.9656 (0.9645 – 0.9667) 0.9850 (0.9838 – 0.9862) 0.6276 (0.6198 – 0.6354) 0.6137 (0.6059 – 0.6215) 0.9921 (0.9902 – 0.9940)
+ Ours 0.9612 (0.9603 – 0.9621) 0.9661 (0.9650 – 0.9672) 0.9858 (0.9847 – 0.9869) 0.6268 (0.6191 – 0.6345) 0.6151 (0.6075 – 0.6227) 0.9938 (0.9919 – 0.9957)
Wan2.1 Vanilla 0.9562 (0.9553 – 0.9571) 0.9656 (0.9644 – 0.9668) 0.9850 (0.9839 – 0.9861) 0.6276 (0.6201 – 0.6351) 0.6137 (0.6061 – 0.6213) 0.9943 (0.9924 – 0.9962)
+ Ours 0.9612 (0.9601 – 0.9623) 0.9661 (0.9649 – 0.9673) 0.9858 (0.9846 – 0.9870) 0.6268 (0.6190 – 0.6346) 0.6151 (0.6073 – 0.6229) 0.9956 (0.9935 – 0.9977)
Appendix J Statistical Validation of Results
Our VBench evaluation uses 4,750 videos (five samples per prompt) across 17 fine-grained dimensions, where even small absolute gains are meaningful. To assess statistical significance, we computed 95% confidence intervals using 10 independent runs, each based on random subsets of 100 videos. As shown in Tables 17, 18, and 19, our method consistently improves or maintains performance across quality-, semantic-, and motion-related metrics without exhibiting bias toward any particular subset. These statistical results provide strong evidence of the robustness and reliability of our approach.
Appendix K User Study Protocol
As presented in Fig. 5 (a), we conduct a human evaluation study to complement automated metrics and directly assess whether ANSE improves video quality and prompt coherence. We evaluate four representative video diffusion models—CogVideoX-2B, CogVideoX-5B, Wan 2.1, and HunyuanVideo—by comparing their outputs with and without ANSE. To construct a fair evaluation protocol, we randomly sample 30 prompts from the Subject Consistency and Overall Consistency dimensions of the VBench benchmark. For each prompt, we generate paired videos using identical noise seeds for both the baseline and the ANSE-equipped models.
We recruit 12 independent human evaluators, all of whom remain fully blind to the models and are restricted to participating only once. Each evaluator is shown two videos side-by-side (baseline vs. ANSE) and responds to two questions:
  1. 1.
    Video Quality: “Which video appears more stable and visually pleasing?”
  2. 2.
    Prompt Alignment: “Which video better represents the given text prompt?”
The order of prompts and the arrangement of video pairs are fully randomized. Across all backbones and prompts, models equipped with ANSE are consistently preferred in terms of both video quality and prompt alignment, demonstrating the practical effectiveness and robustness of ANSE.
Appendix L Additional Qualitative Comparison
@@ Appendix I Effect of Equivalent Computation Budget
Figures 9,  10, 11  12, and  13 present additional examples generated using our noise selection framework including diverse backbone such as, AnimateDiff, CogVideoX2B and 5B, HunyangVideo, and Wan2.1. Across diverse prompts, the selected seeds yield improved spatial detail, aesthetic quality, and semantic alignment, further validating the robustness of our approach. These examples complement our quantitative findings by illustrating the visual impact of BANSA-based noise selection.
@@
Figure 9: Effect of ANSE on semantic fidelity and motion stability in AnimateDiff and FreqPrior. Each block compares baseline generations with those using ANSE-selected noise. In addition, while FreqPrior serves as a noise-refinement baseline, our method is fully compatible and achieves further improvements when combined.
Figure 10: Effect of ANSE on semantic fidelity and motion stability in CogVideoX outputs. Each block compares baseline generations with those using ANSE-selected noise. Across both CogVideoX-2B and 5B, ANSE improves semantic alignment to the prompt and reduces artifacts such as temporal flickering and object distortion. White numbers show BANSA scores for the same prompt.
Figure 11: Additional qualitative comparison of CogVideoX variants with and without ANSE. Results from CogVideoX-2B are shown in the first two rows; the rest show CogVideoX-5B. With ANSE, videos exhibit improved visual quality, better text alignment, and smoother motion transitions compared to the baseline. White numbers show BANSA scores for the same prompt.
Figure 12: Additional qualitative comparison of HunyangVideo with and without ANSE. With ANSE, the generated videos achieve sharper visual fidelity, stronger alignment between text and content, and smoother motion progression than the baseline. White numbers show BANSA scores for the same prompt.
Figure 13: Additional qualitative comparison of Wan2.1 with and without ANSE. ANSE enables videos to deliver higher visual quality, more accurate adherence to the given text, and more seamless motion dynamics compared to the baseline. White numbers show BANSA scores for the same prompt.